# Sparse Hanson-Wright Inequalities with Applications

### v2

@@ Appendix D Proof of Proposition 1
We now prove Proposition 1, following closely the proof of Proposition 4.5 from [24]. We outline the main steps below. First, we recall the stability estimate from Lemma 4.4 of [24]: For $z\in\mathbf{S}$, let $m,\tilde{m}$ be the solutions of the equation $x^{2}+zx+1=0$. If $s$ satisfies $s^{2}+zs+1=r$, then
$\min\{|s-m|,|s-\tilde{m}|\}\leq\sqrt{|r|}.$ (D.1)
@@
$\max\left\{\max_{i\neq j}\|\phi G_{ij}(z_{k})\|_{L_{r}},\max_{i,j\neq k}\|\phi(G_{ij}-G_{ij}^{(k)})(z_{k})\|_{L_{r}},\max_{i}\|\phi Y_{i}G_{ii}(z_{k})\|_{L_{r}}\right\}\leq\zeta_{k}^{2}.$ (D.2)
@@
For $k=0$, we have $\eta_{0}=1$ and $\Gamma(z_{0})\leq 1$ using the trivial bound $\|G(z_{0})\|\leq 1$. Hence, $\phi(z_{0})=1$ and $\mathbb{P}(\Xi_{0})=1$. From the $1+zs+s^{2}=\frac{1}{n}\sum_{i}G_{ii}Y_{i}$ and (D.2), we obtain
@@
It follows from (D.1) that
$\mathbb{P}\left(\min\{|s(z_{0})-m(z_{0})|,|s(z_{0})-\tilde{m}(z_{0})|\}>\zeta_{0}t\right)\leq t^{-2r}.$ (D.3)
@@
$\mathbb{P}(\Omega_{0}^{c})=\mathbb{P}\left(|s(z_{0})-m(z_{0})|>\zeta_{0}t\right)\leq t^{-2r}.$ (D.4)
Likewise, (D.2), together with Markov’s inequality, yields
$\max_{i}\mathbb{P}(|Y_{i}(z_{0})G_{ii}(z_{0})|>\zeta_{0}t)\leq t^{-2r}\quad\text{and}\quad\max_{i\neq j}\mathbb{P}(|G_{ij}(z_{0})|>\zeta_{0}t)\leq t^{-2r}.$ (D.5)
@@
$\mathbb{P}\left(\max_{i,j}|G_{ij}(z_{0})-m(z_{0})\delta_{ij}|>2\zeta_{0}t\right)\leq(n^{2}+2n)t^{-2r}<3n^{2}t^{-2r}.$ (D.7)
@@
Case I. For $1\leq k<\widetilde{\mathcal{K}}$, since $\Gamma(z)$ is $n^{2}$-Lipschitz, we have $\phi(z_{k})=1$ on $\Xi_{k-1}$. Following the derivation of (D.3), we obtain
@@
Similarly to (D.5) and (D.6), the following estimates hold:
$\max_{i}\mathbb{P}(\Omega_{k-1}\cap\Xi_{k-1}\cap\{|G_{ii}(z_{k})-m(z_{k})|>2\zeta_{k}t\})\leq 2t^{-2r},$ (D.8)
$\max_{i\neq j}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ij}(z_{k})|>\zeta_{k}t\})\leq t^{-2r},$ (D.9)
$\max_{i,j\neq l}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ij}(z_{k})-G_{ij}^{(l)}(z_{k})|>\zeta_{k}t\})\leq t^{-2r}.$ (D.10)
@@
Finally, similar to (D.7), we arrive at
@@
$\mathbb{P}(\Xi_{k-1}\cap\{|s(z_{k})-m(z_{k})|>6\zeta_{k}t)\leq t^{-2r},$ (D.11)
$\max_{i}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ii}(z_{k})-m(z_{k})|>10\zeta_{k}t\})\leq 2t^{-2r},$ (D.12)
$\max_{i\neq j}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ij}(z_{k})|>\zeta_{k}t\})\leq t^{-2r}.$ (D.13)
@@
Finally, following an argument similar to (D.7), we conclude

### v3

@@ B.2 Proof of Theorem 9
We now prove Proposition 7.1, following closely the proof of Proposition 4.5 from [34]. We outline the main steps below. First, we recall the stability estimate from Lemma 4.4 of [34]: For $z\in\mathbf{S}$, let $m,\tilde{m}$ be the solutions of the equation $x^{2}+zx+1=0$. If $s$ satisfies $s^{2}+zs+1=r$, then
$\min\{|s-m|,|s-\tilde{m}|\}\leq\sqrt{|r|}.$ (C.1)
@@
$\max\left\{\max_{i\neq j}\|\phi G_{ij}(z_{k})\|_{L_{r}},\max_{i,j\neq k}\|\phi(G_{ij}-G_{ij}^{(k)})(z_{k})\|_{L_{r}},\max_{i}\|\phi Y_{i}G_{ii}(z_{k})\|_{L_{r}}\right\}\leq\zeta_{k}^{2}.$ (C.2)
@@
For $k=0$, we have $\eta_{0}=1$ and $\Gamma(z_{0})\leq 1$ using the trivial bound $\|G(z_{0})\|\leq 1$. Hence, $\phi(z_{0})=1$ and $\mathbb{P}(\Xi_{0})=1$. From the $1+zs+s^{2}=\frac{1}{n}\sum_{i}G_{ii}Y_{i}$ and (C.2), we obtain
@@
It follows from (C.1) that
$\mathbb{P}\left(\min\{|s(z_{0})-m(z_{0})|,|s(z_{0})-\tilde{m}(z_{0})|\}>\zeta_{0}t\right)\leq t^{-2r}.$ (C.3)
@@
$\mathbb{P}(\Omega_{0}^{c})=\mathbb{P}\left(|s(z_{0})-m(z_{0})|>\zeta_{0}t\right)\leq t^{-2r}.$ (C.4)
Likewise, (C.2), together with Markov’s inequality, yields
$\max_{i}\mathbb{P}(|Y_{i}(z_{0})G_{ii}(z_{0})|>\zeta_{0}t)\leq t^{-2r}\quad\text{and}\quad\max_{i\neq j}\mathbb{P}(|G_{ij}(z_{0})|>\zeta_{0}t)\leq t^{-2r}.$ (C.5)
@@
$\mathbb{P}\left(\max_{i,j}|G_{ij}(z_{0})-m(z_{0})\delta_{ij}|>2\zeta_{0}t\right)\leq(n^{2}+2n)t^{-2r}<3n^{2}t^{-2r}.$ (C.7)
@@
Case I. For $1\leq k<\widetilde{\mathcal{K}}$, since $\Gamma(z)$ is $n^{2}$-Lipschitz, we have $\phi(z_{k})=1$ on $\Xi_{k-1}$. Following the derivation of (C.3), we obtain
@@
Similarly to (C.5) and (C.6), the following estimates hold:
$\max_{i}\mathbb{P}(\Omega_{k-1}\cap\Xi_{k-1}\cap\{|G_{ii}(z_{k})-m(z_{k})|>2\zeta_{k}t\})\leq 2t^{-2r},$ (C.8)
$\max_{i\neq j}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ij}(z_{k})|>\zeta_{k}t\})\leq t^{-2r},$ (C.9)
$\max_{i,j\neq l}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ij}(z_{k})-G_{ij}^{(l)}(z_{k})|>\zeta_{k}t\})\leq t^{-2r}.$ (C.10)
@@
Finally, similar to (C.7), we arrive at
@@
$\mathbb{P}(\Xi_{k-1}\cap\{|s(z_{k})-m(z_{k})|>6\zeta_{k}t)\leq t^{-2r},$ (C.11)
$\max_{i}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ii}(z_{k})-m(z_{k})|>10\zeta_{k}t\})\leq 2t^{-2r},$ (C.12)
$\max_{i\neq j}\mathbb{P}(\Xi_{k-1}\cap\{|G_{ij}(z_{k})|>\zeta_{k}t\})\leq t^{-2r}.$ (C.13)
@@
Finally, following an argument similar to (C.7), we conclude
@@
Appendix D Proofs of Corollary 1 and Corollary 2
D.1 Proof of Corollary 1
To obtain the Hanson-Wright inequality for a general matrix $A$, we bound the concentration of the diagonal sum $S_{\mathrm{diag}}=\sum_{i=1}a_{ii}X_{i}^{2}$ in the quadratic form $\sum_{i,j}a_{ij}X_{i}X_{j}$ by applying Theorem 2 to $X_{i}^{2}-EX_{i}^{2}$. Note that
$\mathbb{E}|X_{i}^{2}-\mathbb{E}X_{i}^{2}|^{l}\leq 2^{l}\mathbb{E}|X_{i}|^{2l}\leq 2^{l}p_{i}K^{2l}(2l)^{\frac{2l}{\alpha}}.$ (D.1)
Hence, $X_{i}^{2}-EX_{i}^{2}$’s are sparse $\frac{1}{2}$-subexponential random variables. In particular, the tail bound for $S_{\mathrm{diag}}$ is given by
$\mathbb{P}\left(\left|S_{\mathrm{diag}}-\mathbb{E}S_{\mathrm{diag}}\right|\geq t\right)\leq\,e^{2}\exp\left(-c\min\left\{\frac{t^{2}}{K^{4}\sum_{i}a_{ii}^{2}p_{i}},\left(\frac{t}{K^{2}\max_{i}|a_{ii}|}\right)^{\frac{1}{2}}\right\}\right).$ (D.2)
Combining the bounds for the diagonal and off-diagonal terms, we prove Corollary 1.
D.2 Proof of Corollary 2
In this section, we briefly discuss the extension of our results to the concentration of quadratic forms of noncentered random variables. Let $Y_{1},\cdots,Y_{n}$ be independent random variables, and consider a symmetric, diagonal-free matrix $A$ for simplicity. Denote $X_{i}=Y_{i}-\mathbb{E}(Y_{i})$. We rewrite
$Y^{T}AY=\sum_{i\neq j}a_{ij}Y_{i}Y_{j}=\sum_{i\neq j}a_{ij}(X_{i}+\mathbb{E}(Y_{i}))(X_{j}+\mathbb{E}(Y_{j})).$ (D.3)
Through straightforward calculation, we obtain
The concentration inequality for $Y^{T}AY-\mathbb{E}(Y^{T}AY)$ can be obtained by considering those of $X^{T}AX$ and $\sum_{i}b_{i}X_{i}$ separately. Assume the moments of $Y_{i}$ satisfy $\mathbb{E}|Y_{i}|^{r}\leq p_{i}(Kr)^{r}$ for $r\geq 1$. Then from the inequality $\|X_{i}\|_{L_{r}}=\|Y_{i}-\mathbb{E}(Y_{i})\|_{L_{r}}\leq 2\|Y_{i}\|_{L_{r}}$, we deduce that the $X_{i}$’s are independent centered random variables satisfying $\mathbb{E}|X_{i}|^{r}\leq p_{i}(2Kr)^{r}$ for $r\geq 2$. Consequently, the conclusion of Theorem 1 holds for $X^{T}AX$ where $\alpha=1$. Moreover, $\sum_{i=1}^{n}b_{i}X_{i}$ is a sum of independent random variables. Applying Theorem 2 with $\alpha=1$, we find that the tail bound for $\sum_{i=1}^{n}b_{i}X_{i}$ is a mixture of sub-gaussian and subexponential tails. For the subexponential tail, we can use the estimate that
The conclusion of Corollary 2 follows by combining the results of Theorem 1 and Theorem 2 in the preceding discussion.